Problem 3: Calculating a portfolios beta and CAPM-based expected rate of return Ashley is curious to know what her portfolios CAPM-based expected rate of return should be. After doing some research, she determines that the current market values and betas of each of her 5 stock are as listed below. She is informed by her financial advisor that the risk-free rate is 3% and the market risk premium is 8%. Calculate the expected rate of return on Ashleys portfolio. Stock Market Value Beta $35,000 $40,000 $45,000 $50,000 $80,000 1.6 1.2 1.0 0.8 o.8
Answer #1

Please provide rating.............

Expected rerurn on stock = Risk free rate + Market risk premium * beta
i ii iii iv v=iii+(iv)*iii vi=i*v
Stock Market value Beta Risk free rate Market risk premium Expected return Expected return * value
A 35000 1.6 3% 8% 15.80% 5530
B 40000 1.2 3% 8% 12.60% 5040
C 45000 1 3% 8% 11.00% 4950
D 50000 -0.8 3% 8% -3.40% -1700
E 80000 0.8 3% 8% 9.40% 7520
250000 21340
Expected return on portfolio = 21340/250000 8.54%